OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form

OpenStudy (aroub):

y=mx+b , m=-1, (-4,2) 2=-1(-4)+b 2=4+b -2=b =>y=-x-2

OpenStudy (lgbasallote):

aww you ran out of those one variable graphs :( bummer...those were easier -___- haha

OpenStudy (anonymous):

Anybody interested in a trick?

OpenStudy (aroub):

Yup!

OpenStudy (lgbasallote):

does this involve making someting disappear *_* hahaha kidding...go on...

OpenStudy (anonymous):

lol

OpenStudy (aroub):

lol yeah! That will be fun :)

OpenStudy (anonymous):

Any line paralle to \( ax + by =c \) will be for the form \( ax+by=k \), now here we have \( x+y=0 (c) \) substituting we will get \(-4+2 =2 \) so the equation will be \( x+y=-2\)

OpenStudy (aroub):

Oh that's cool :D

OpenStudy (anonymous):

:)